OpenStudy (anonymous):

Medal and follow.... Choose the equation of the horizontal line that passes through the point (-8, -7)

OpenStudy (anonymous):

@Mandre

OpenStudy (anonymous):

Plz

OpenStudy (anonymous):

@shy

OpenStudy (anonymous):

@shyyy!

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

I need help

hartnn (hartnn):

equation of horizontal line is of the form y = c where 'c' is the y co-ordinate of any point on the line

hartnn (hartnn):

what is the y co-ordinate here?

OpenStudy (anonymous):

-7

OpenStudy (anonymous):

im sorry but i dont remember how to do this

OpenStudy (anonymous):

The two are (-8,-7)

hartnn (hartnn):

-7 is the y co-ordinate, correct! so the equation is just y = -7 thats it! :)

OpenStudy (anonymous):

Wow that was easy Can u help me with 1 more easy one

hartnn (hartnn):

sure :)

OpenStudy (anonymous):

Ok which of the following equations best represent the graph of f(n)

OpenStudy (anonymous):

A f(y)=n B f(x)=n C x=f(n) D y=f(n)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@har

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

where is the graph ?

OpenStudy (anonymous):

There is none

hartnn (hartnn):

then the question is incomplete!

hartnn (hartnn):

best represent the \(\huge graph\) of f(n) which graph ?

OpenStudy (anonymous):

No i coped it word by word look at the question carefully and the answers

hartnn (hartnn):

ok

hartnn (hartnn):

graph of f(x) mean y = f(x) so graph of f(n) would be y = f(n)

OpenStudy (anonymous):

REPRESENT the graph of f(n) i think it means if there were a graph how would u write the answer

OpenStudy (anonymous):

Oh thx soo much

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

Ok